# Harmonic $1$-forms on real loci of Calabi-Yau manifolds

### v1

@@ 5.1 Approximate Calabi-Yau metrics
Apart from this, it is folklore that for Calabi-Yau manifold that is a smoothing of a singular Calabi-Yau manifold, the approximate metric is only a poor fit near the smoothed singularity, and is a good fit far away from the smoothed singularity. This is confirmed by the data presented in Fig. 1.
Figure 1: Plots showing the Monge–Ampère error $|1-\omega^{m}/\Omega\wedge\overline{\Omega}|$ on the $y$-axis over the approximate distance to a singular locus. From left to right: CICY1, CICY2, then two plots for Quintic. The approximate distance to a singular locus is computed as follows: for CICY1 and CICY2 as $\min\{d_{2}(x)/|x|^{2},\,d_{4}(x)/|x|^{4}\}$, where $d_{2}(x)=\max\{|x_{4}^{2}+x_{5}^{2}|,\,|x_{0}|^{2},\,|x_{1}|^{2},\,|x_{2}|^{2},\,|x_{3}|^{2}\}$ and $d_{4}(x)=\max\{|x_{1}^{4}+x_{2}^{4}+x_{3}^{4}|/|x|^{4},\,|x_{0}|^{4},\,|x_{4}|^{4},\,|x_{5}|^{4}\}$. Quintic is a smoothing of a variety with two singularities, namely (a) the intersection of the divisors $Z(f)$ and $Z(g^{*})$ from Section 3.2.2, and (b) an ordinary double point at $a=[1:0:0:0:0]$. The $x$-axis of the third image is an approximation of the distance to (a), namely $\max\{|x_{0}^{2}+\dots+x_{4}^{2}|/|x|^{2},\,f^{*}(x)/|x|^{3}\}$. The $x$-axis of the fourth image is an approximation of the distance to (b), namely $|x-(|x|,0,0,0,0)|/|x|$.
@@
Figure 2: Plots showing the formation of long necks. From left to right: Quintic, CICY1, CICY2. The three varieties are perturbations of singular varieties. The horizontal axis measures an approximation of the distance to the singular locus of the initial singular variety, namely $\max\{|x_{0}^{2}+\dots+x_{4}^{2}|/|x|^{2},\,f^{*}(x)/|x|^{3}\}$ for Quintic and $\min\{d_{2}(x)/|x|^{2},\,d_{4}(x)/|x|^{4}\}$, where $d_{2}(x)=\max\{|x_{4}^{2}+x_{5}^{2}|,\,|x_{0}|^{2},\,|x_{1}|^{2},\,|x_{2}|^{2},\,|x_{3}|^{2}\}$ and $d_{4}(x)=\max\{|x_{1}^{4}+x_{2}^{4}+x_{3}^{4}|/|x|^{4},\,|x_{0}|^{4},\,|x_{4}|^{4},\,|x_{5}|^{4}\}$ for CICY1 and CICY2. The vertical axis shows $\max_{v\in T_{x}M:|v|_{FS}=1}|v|_{h}$, where $|\cdot|_{FS}$ denotes the length of a tangent vector in the ambient Fubini-Study metric and $|\cdot|_{h}$ denotes the length of a vector in the learned approximate Calabi-Yau metric. For Quintic and CICY2, which are small perturbations of a singular variety, the data shows a long neck forming. CICY1 is a large perturbation of a singular variety and the neck formation cannot be observed.
@@ 5.2 Approximate harmonic $1$-forms
  • The real locus of Quintic does admit a non-zero harmonic $1$-form, but it must have zeros by Proposition 3.10.
@@
Table 2: Results of approximately computing a harmonic $1$-form on four example varieties. Harmonic loss measures the quality of the approximate solution. On varieties whose real locus has zero fist Betti number, we expect this quality to be low. Normalised min norm is the minimum norm in a large number of sampled points of the approximate $1$-form scaled to have unit $L^{2}$-norm. On some varieties, we expect the harmonic $1$-form to have no zeros, on others we have shown it must have zeros. This is explained in Section 5.2.
Figure 3: Plot of the data from Table 2. $x$-axis shows the approximation error of the approximate $1$-form $\omega$, $y$-axis shows the minimum value of $|\omega|$ over the manifold. The data confirms our expectation and neatly fits into a pattern of large/small approximation error and having/not having a zero.
@@
The data in Figs. 3 and 2 suggest that there exists a harmonic $1$-form on the real locus $L\cong S^{1}\times S^{2}$ of CICY1. Inspection of the approximate solution shows that it is approximately zero in the $S^{1}$-direction, and approximately a $1$-form flowing from a single source to a single sink in the $z_{2}$-direction on $S^{2}$. This is visualised in Fig. 5. We note that such a $1$-form on $S^{2}$ would be approximately the differential of an eigenfunction of the Laplacian with minimum eigenvalue, so it minimises the functional $(\left|\left|\operatorname{d}\!\omega\right|\right|_{L^{2}}+\left|\left|\operatorname{d}\!^{*}\omega\right|\right|_{L^{2}})/\left|\left|\omega\right|\right|_{L^{2}}$ among forms on $S^{2}$. This result is somewhat unexpected, but we found that the neural network learned this solution, even when initialised to be the solution from CICY2, which is far from non-zero in the $S^{1}$-direction.
Figure 5: First three pictures show the approximate harmonic $1$-form restricted to different one-dimensional submanifolds of CICY1. Fourth picture shows the $1$-form on the variety $-1+z_{1}^{4}+z_{2}^{4}+z_{3}^{4}=0$. (Note that this variety is not contained in CICY1, but only nearby.) For better visualisation, arrow lengths have been rescaled as follows: $100$ (top-left), $0.5$ (remaining three pictures). The plots show that the $1$-form is approximately zero in the $S^{1}$-direction, and on $S^{2}$ appears to have a single sink and a single source.

### v2

@@ 5.1 Approximate Calabi-Yau metrics
Apart from this, it is folklore that for a Calabi-Yau manifold that is a smoothing of a singular Calabi-Yau manifold, the approximate metric is only a poor fit near the smoothed singularity, and is a good fit far away from the smoothed singularity. This is confirmed by the data presented in Fig. 1.
Figure 1: Plots showing the Monge–Ampère error $|1-\omega^{m}/\Omega\wedge\overline{\Omega}|$ on the $y$-axis over the approximate distance to a singular locus. From left to right: CICY1, CICY2, then two plots for Quintic. The approximate distance to a singular locus is computed as follows: for CICY1 and CICY2 as $\min\{d_{2}(x)/|x|^{2},\,d_{4}(x)/|x|^{4}\}$, where $d_{2}(x)=\max\{|x_{4}^{2}+x_{5}^{2}|,\,|x_{0}|^{2},\,|x_{1}|^{2},\,|x_{2}|^{2},\,|x_{3}|^{2}\}$ and $d_{4}(x)=\max\{|x_{1}^{4}+x_{2}^{4}+x_{3}^{4}|,\,|x_{0}|^{4},\,|x_{4}|^{4},\,|x_{5}|^{4}\}$. Quintic is a smoothing of a variety with two singularities, namely (a) the intersection of the divisors $Z(f)$ and $Z(g^{*})$ from Section 3.2.2, and (b) an ordinary double point at $a=[1:0:0:0:0]$. The $x$-axis of the third image is an approximation of the distance to (a), namely $\max\{|x_{0}^{2}+\dots+x_{4}^{2}|/|x|^{2},\,f^{*}(x)/|x|^{3}\}$. The $x$-axis of the fourth image is an approximation of the distance to (b), namely |x(|x|,0,0,0,0)|/|x||x-(|x|,0,0,0,0)|/|x|.
@@
Figure 2: Plots showing the formation of long necks. From left to right: Quintic, CICY1, CICY2. The three varieties are perturbations of singular varieties. The horizontal axis measures an approximation of the distance to the singular locus of the initial singular variety, namely $\max\{|x_{0}^{2}+\dots+x_{4}^{2}|/|x|^{2},\,f^{*}(x)/|x|^{3}\}$ for Quintic and $\min\{d_{2}(x)/|x|^{2},\,d_{4}(x)/|x|^{4}\}$, where $d_{2}(x)=\max\{|x_{4}^{2}+x_{5}^{2}|,\,|x_{0}|^{2},\,|x_{1}|^{2},\,|x_{2}|^{2},\,|x_{3}|^{2}\}$ and $d_{4}(x)=\max\{|x_{1}^{4}+x_{2}^{4}+x_{3}^{4}|,\,|x_{0}|^{4},\,|x_{4}|^{4},\,|x_{5}|^{4}\}$ for CICY1 and CICY2. The vertical axis shows $\max_{v\in T_{x}M:|v|_{FS}=1}|v|_{h}$, where $|\cdot|_{FS}$ denotes the length of a tangent vector in the ambient Fubini-Study metric and $|\cdot|_{h}$ denotes the length of a vector in the learned approximate Calabi-Yau metric. For Quintic and CICY2, which are small perturbations of a singular variety, the data shows a long neck forming. CICY1 is a large perturbation of a singular variety and the neck formation cannot be observed.
@@ 5.2 Approximate harmonic $1$-forms
  • The real locus of Quintic does admit a non-zero harmonic $1$-form, but it must have at least one zero by Proposition 3.10.
@@
Table 2: Results of approximately computing a harmonic $1$-form on four example varieties. Harmonic loss measures the quality of the approximate solution. On varieties whose real locus has zero first Betti number, we expect this quality to be low. Normalised min norm is the minimum norm in a large number of sampled points of the approximate 11-form scaled to have unit $L^{2}$-norm. On some varieties, we expect the harmonic $1$-form to have no zeros, on others we have shown it must have zeros. This is explained in Section 5.2.
Figure 3: Plot of the data from Table 2. $x$-axis shows the approximation error of the approximate 11-form $\omega$, $y$-axis shows the minimum value of $|\omega|$ over the manifold. The data confirms our expectation and neatly fits into a pattern of large/small approximation error and having/not having a zero.
@@
[…]
Figure 5: First three pictures show the approximate harmonic $1$-form restricted to different one-dimensional submanifolds of CICY1. Fourth picture shows the $1$-form on the variety $1+z_{1}^{4}+z_{2}^{4}+z_{3}^{4}=0$. (Note that this variety is not contained in CICY1, but only nearby.) For better visualisation, arrow lengths have been rescaled as follows: $100$ (top-left), $0.5$ (remaining three pictures). The plots show that the $1$-form is approximately zero in the $S^{1}$-direction, and on $S^{2}$ appears to have a single sink and a single source.